# ELECTRA and GPT-4o: Cost-Effective Partners for Sentiment Analysis

### v1

@@ Abstract
Bidirectional transformers excel at sentiment analysis, and Large Language Models (LLM) are effective zero-shot learners. Might they perform better as a team? This paper explores collaborative approaches between ELECTRA and GPT-4o for three-way sentiment classification. We fine-tuned (FT) four models (ELECTRA Base/Large, GPT-4o/4o-mini) using a mix of reviews from Stanford Sentiment Treebank (SST) and DynaSent. We provided input from ELECTRA to GPT as: predicted label, probabilities, and retrieved examples. Sharing ELECTRA Base FT predictions with GPT-4o-mini significantly improved performance over either model alone (82.74 macro F1 vs. 79.29 ELECTRA Base FT, 79.52 GPT-4o-mini) and yielded the lowest cost/performance ratio (\$0.12/F1 point). However, when GPT models were fine-tuned, including predictions decreased performance. GPT-4o FT-M was the top performer (86.99), with GPT-4o-mini FT close behind (86.77) at much less cost (\$0.38 vs. \$1.59/F1 point). Our results show that augmenting prompts with predictions from fine-tuned encoders is an efficient way to boost performance, and a fine-tuned GPT-4o-mini is nearly as good as GPT-4o FT at 76% less cost. Both are affordable options for projects with limited resources.
@@ 1 Introduction
Recent work has explored collaboration between these models, such as using GPT to augment data of minority classes before fine-tuning with RoBERTa (Kok-Shun et al., 2023), using GPT for aspect extraction and RoBERTa for sentiment scoring (Qian et al., 2024), and escalating to LLMs when RoBERTa classification confidence was low (Andrade et al., 2024). However, leveraging predictions from fine-tuned encoders to enhance LLMs remains under-explored.
@@
To test these hypotheses, we established four baselines and conducted 23 experiments across three sentiment classification datasets: Stanford Sentiment Treebank (SST), and DynaSent Rounds 1 and 2. We used ELECTRA Base/Large and GPT-4o/4o-mini, each of which were fine-tuned (FT) on a merge of SST and DynaSent reviews.
We investigated the effects of different prompt augmentation scenarios using DSPy (Khattab et al., 2024): sharing the predicted class text label, the probabilities of each class as percentages, retrieving similar reviews with their class labels, and combinations. We evaluated classification performance with the macro average F1 score, and cost-effectiveness by dividing the total fine-tuning costs by the F1. Our key insights are the following.
@@
Adding probabilities or examples helped. Using probabilities, or including few-shot examples, both improved performance for GPT-4o (not fine-tuned) vs. the predicted label alone. These only helped GPT-4o-mini on DynaSent R2 and SST-3.
@@
Sharing predictions hurt fine-tuned GPTs. When GPT models were fine-tuned, including ELECTRA predictions decreased performance—even when fine-tuned with the same inference-time prompt that included the ELECTRA prediction.
@@ 2.1 MLMs and ELECTRA
ELECTRA was found to be a top performer in sentiment classification on datasets such as SST (Clark et al., 2020), DynaSent (Potts et al., 2021), and IMDB movie reviews (B et al., 2023). It’s for these reasons that we chose to use ELECTRA in our research, in addition to observing a performance gain relative to RoBERTa in early trials.
@@ 3 Data
Models were trained and evaluated on a merge of movie reviews from the Stanford Sentiment Treebank (SST) (Socher et al., 2013) and business reviews from DynaSent Rounds 1 and 2 (Potts et al., 2021). See Table 1 for examples. By default, SST is a five-way classification (positive, somewhat positive, neutral, somewhat negative, negative). The positive and negative classes were combined to produce SST-3 (positive, neutral, negative).
The SST-3, DynaSent R1, and DynaSent R2 datasets were randomly mixed to form a Merged dataset with 102,097 Train examples, 5,421 Validation examples, and 6,530 Test examples. See Table 2 for the distribution of labels, and Table 3 for a breakdown of sources.
It’s worth noting that the source datasets all have class imbalances. Merging the data helps mitigate this imbalance, but there is still a majority of neutral examples in the training split. Another potential issue is that the models will learn the dominant dataset, which is DynaSent R1. As a test, the minority classes were over-sampled to create a new balanced dataset. When this was evaluated, the performance did not improve.
@@ 4 Models
ELECTRA (Clark et al., 2020) was chosen as the bidirectional transformer because its pre-training architecture gives it an advantage over MLMs. It also outperformed RoBERTa in early trials. We evaluated both the Base (110M parameters) and Large (335M parameters) variants.
To function as a classifier, ELECTRA’s output is sent through a mean pooling layer. A classifier head is appended with 2 hidden layers of dimension 1024, and a final output dimension of 3. Swish GLU (Shazeer, 2020) was used as the hidden activation function, and dropout layers were added with a rate of 0.3. See Appendix A for more details on the model architecture and hyper-parameters.
@@ 5.1 ELECTRA Baseline & Fine-tuning
Two baseline models were established by training only classifier heads for ELECTRA Base and Large. Hyper-parameters were consistent with the fully fine-tuned versions. The fine-tuning process involved a number of trials on Lambda Labs multi-GPU instances to identify the best hyper-parameters, optimizer, and learning rate schedule. See Appendix A for the final configuration.
@@ 5.2 GPT Data Preparation & Fine-tuning
To use OpenAI’s fine-tuning API, we converted the Merged training data to JSONL format that defined the System, User, and Assistant roles. We noticed that if the context at inference time varied even slightly from the fine-tuning context, performance would suffer. So we created three templates to enable better comparisons between fine-tuned and default models using the same DSPy signatures (see Appendix B):
  • Minimal (FT-M): No prompt other than System role.
@@
  • Prompt with Label (FT-L: User role included DSPy prompt with ELECTRA predicted label.
We included the ELECTRA predictions in the third template to align the fine-tuning context with the inference time context, but also to provide an opportunity for the GPT models to learn from the ELECTRA predictions. In total there were 6 fine-tuning jobs (see Table 5).
@@
Model Code Format Tokens
GPT-4o-mini FT Prompt 11.0M
GPT-4o-mini FT-L Prompt w/Base Label 13.2M
GPT-4o-mini FT-M Minimal 5.5M
GPT-4o FT Prompt 11.0M
GPT-4o FT-L Prompt w/Large Label 13.2M
GPT-4o FT-M Minimal 5.5M
@@ 5.3 DSPy Signatures & Modules
Using DSPy, we explored a variety of approaches to integrating ELECTRA’s output into GPT’s decision-making process. Each approach was implemented as a custom DSPy signature and module (see Appendix C for the full examples).
@@
Balanced Examples. Similar to Top Examples, except a different retriever was used that retrieved a total of six examples (two from each class) to ensure the few-shot examples with true labels did not bias the answer toward a particular class—although that might be desirable.
All of the Above. And lastly, a final DSPy signature had all of the above context from ELECTRA included: classification prompt, predicted label, probabilities, and top five examples—not balanced.
We then conducted two of the four baselines, and 21 of the 23 experiments (see Table 6) using these DSPy signatures and modules. The fine-tuned ELECTRA models and retriever were instantiated locally for inference, and the GPT models were accessed via OpenaAI API.
@@ 6 Results
Table 6: Summary of Model Configurations, Performance, and Cost
Merged2 DynaSent R1 DynaSent R2 SST-3 Cost (\$)3

### v2

@@ Abstract
Bidirectional transformers excel at sentiment analysis, and Large Language Models (LLM) are effective zero-shot learners. Might they perform better as a team? This paper explores collaborative approaches between ELECTRA and GPT-4o for three-way sentiment classification. We fine-tuned (FT) four models (ELECTRA Base/Large, GPT-4o/4o-mini) using a mix of reviews from Stanford Sentiment Treebank (SST) and DynaSent. We provided input from ELECTRA to GPT as: predicted label, probabilities, and retrieved examples. Sharing ELECTRA Base FT predictions with GPT-4o-mini significantly improved performance over either model alone (82.50 macro F1 vs. 79.14 ELECTRA Base FT, 79.41 GPT-4o-mini) and yielded the lowest cost/performance ratio (\$0.12/F1 point). However, when GPT models were fine-tuned, including predictions decreased performance. GPT-4o FT-M was the top performer (86.99), with GPT-4o-mini FT close behind (86.70) at much less cost (\$0.38 vs. \$1.59/F1 point). Our results show that augmenting prompts with predictions from fine-tuned encoders is an efficient way to boost performance, and a fine-tuned GPT-4o-mini is nearly as good as GPT-4o FT at 76% less cost. Both are affordable options for projects with limited resources.
@@ 1 Introduction
Recent work has explored collaboration between these models, such as using GPT to augment data of minority classes before fine-tuning with RoBERTa (Kok-Shun et al., 2023), using GPT for aspect extraction and RoBERTa for sentiment scoring (Qian et al., 2024), and escalating to LLMs when RoBERTa classification confidence was low (Andrade et al., 2024). However, leveraging external knowledge of sentiment from fine-tuned encoders to enhance LLMs remains under-explored.
@@
To test these hypotheses, we established four baselines and conducted 28 experiments across three sentiment classification datasets: Stanford Sentiment Treebank (SST), and DynaSent Rounds 1 and 2. We used ELECTRA Base/Large and GPT-4o/4o-mini, each of which were fine-tuned (FT) on a merge of SST and DynaSent reviews.
We investigated the effects of different prompt augmentation scenarios using Declarative Self-improving Python (DSPy) (Khattab et al., 2024), a framework for programming language models. We started with a prompt to classify a review, and augmented it with knowledge from ELECTRA in the form of: the predicted class label, the probabilities of each class, similar reviews with their class labels, and combinations. We evaluated classification performance with the macro average F1 score, and cost-effectiveness by dividing total fine-tuning costs by the F1. Our key insights are the following.
@@
Adding probabilities or examples did not help. Using probabilities, or including few-shot examples, did not improve performance more than the predicted label alone for both GPT models.
@@
Sharing predictions hurt fine-tuned GPTs. When GPT models were fine-tuned, including ELECTRA predictions decreased performance—even when fine-tuned with the same inference-time prompt that included the ELECTRA prediction. Fine-tuning with the prediction for more epochs allowed GPT to discriminate better. Performance improved, but the cost grew significantly.
@@ 2.1 MLMs and ELECTRA
ELECTRA was found to be a top performer in sentiment classification on datasets such as SST (Clark et al., 2020), DynaSent (Potts et al., 2021), and IMDB movie reviews (B et al., 2023). It was also found to be better suited for prompt-based learning due to its use of a discrimnator (Xia et al., 2022). We chose to use ELECTRA for these reasons, in addition to observing a performance gain relative to RoBERTa in early trials.
@@ 3 Data
Models were trained and evaluated in English on a merge of movie reviews from the Stanford Sentiment Treebank (SST) (Socher et al., 2013) and business reviews from DynaSent Rounds 1 and 2 (Potts et al., 2021), licensed under Apache 2.0 and Creative Commons Attribution 4.0 respectively. See Table 1 for examples. By default, SST is a five-way classification (positive, somewhat positive, neutral, somewhat negative, negative). The positive and negative classes were combined to produce SST-3 (positive, neutral, negative).
The SST-3, DynaSent R1, and DynaSent R2 datasets were randomly mixed to form a Merged dataset with 102,097 Train examples, 5,421 Validation examples, and 6,530 Test examples. See Table 2 for the distribution of labels, and Table 3 for a breakdown of sources. It’s worth noting that the source datasets all have class imbalances. Merging the data helps mitigate this imbalance, but there is still a majority of neutral examples in the training split. Another potential issue is that the models will learn the dominant dataset, which is DynaSent R1. As a test, the minority classes were over-sampled to create a new balanced dataset. When this was evaluated, the performance did not improve.
@@ 4 Models
ELECTRA (Clark et al., 2020), released with an Apache 2.0 license, was chosen as the bidirectional transformer because its pre-training architecture gives it an advantage over MLMs. It also outperformed RoBERTa in early trials. We evaluated both the Base (110M parameters) and Large (335M parameters) variants.
To function as a classifier, ELECTRA’s output is sent through a mean pooling layer. A classifier head is appended with 2 hidden layers of dimension 1024, and a final output dimension of 3. Swish GLU (Shazeer, 2020) was used as the hidden activation function, and dropout layers were added with a rate of 0.3. See Appendix B for more details on the model architecture and hyper-parameters.
@@ 5.1 ELECTRA Baseline & Fine-tuning
Two baseline models were established by training only classifier heads for ELECTRA Base and Large. Hyper-parameters were consistent with the fully fine-tuned versions. The fine-tuning process involved a number of trials on Lambda Labs multi-GPU instances to identify the best hyper-parameters, optimizer, and learning rate schedule. See Appendix B for the final configuration.
@@ 5.2 GPT Data Preparation & Fine-tuning
To use OpenAI’s fine-tuning API, we converted the Merged training data to JSONL format that defined the System, User, and Assistant roles. We noticed that if the context at inference time varied even slightly from the fine-tuning context, performance would suffer. So we created three templates to enable better comparisons between fine-tuned and default models using the same DSPy signatures (see Appendix D):
  • Minimal (FT-M): No prompt other than System role. User role only contained the review sentence.
@@
  • Prompt with Label (FT-L): User role included DSPy prompt with ELECTRA predicted label.
We included the ELECTRA predictions in the third template to align the fine-tuning context with the inference time context, but also to provide an opportunity for the GPT models to learn from the ELECTRA predictions. In total there were 9 fine-tuning jobs (see Table 5, and Appendix C for GPT fine-tuning details).
@@
Model Code Format Epochs
4o-mini FT-M Minimal 1
4o-mini FT Prompt 1
4o-mini FT-L Prompt w/Base Label 1
4o-mini FT-L Prompt w/Base Label 5
4o-mini FT-L Prompt w/Large Label 1
4o-mini FT-L Prompt w/Large Label 5
4o FT-M Minimal 1
4o FT Prompt 1
4o FT-L Prompt w/Large Label 1
@@ 5.3 DSPy Signatures & Modules
With DSPy, you create modules (ex: Classify, ClassifyWithExamples), signatures (input/output templates, ex: review + examples $\rightarrow$ classification), define metrics (ex: classification_match) and evaluators of data, and use optimizers to find the best performing prompt or module parameters.
We explored a variety of approaches to integrating ELECTRA’s output into GPT’s decision-making process. Each approach was implemented as a custom DSPy signature and module (see Appendix E for the full examples).
@@
Balanced Examples. If ELECTRA was wrong, and the Top Examples were all of the same class, it might be hard for GPT to make an independent decision. To compensate, in Balanced Examples, a different retriever was used that retrieved a total of six examples (the top two examples from each class) to ensure the few-shot examples with true labels did not bias the answer toward a particular class—although that might be desirable sometimes.
All of the Above. And lastly, a final DSPy signature had all of the above context from ELECTRA included: classification prompt, predicted label, probabilities, and top five examples (not balanced). It was unclear if providing all this information would help GPT make a decision, or if the large number of tokens would dilute the signal.
We then conducted two of the four baselines, and 26 of the 28 experiments (see Table 6) using these DSPy signatures and modules. The fine-tuned ELECTRA models and retriever were instantiated locally for inference, and the GPT models were accessed via OpenaAI API. To address single-run concerns, each baseline and experiment was run a second time with a different random seed (123 vs. 42) and temperature (0.1 vs. 0.0).
@@ 6 Results
Table 6: Summary of Model Configuration, Test Set Performance, and Cost
Merged3 DynaSent R1 DynaSent R2 SST-3 Cost (\$)5
ID1 GPT2 ELECTRA Description Macro F14 Macro F14 Macro F14 Macro F14 FT /F1
B1 Base Baseline, Classifier head 69.51 ± 0.20 70.86 ± 0.15 61.39 ± 0.28 60.60 ± 0.36 0.65 0.01
B2 Large Baseline, Classifier head 67.94 ± 0.08 69.70 ± 0.04 59.78 ± 0.00 57.95 ± 0.37 2.51 0.04
B3 4o-mini Baseline (Zero shot) 79.41 ± 0.16 81.16 ± 0.05 77.02 ± 0.47 69.99 ± 0.97
B4 4o Baseline (Zero shot) 79.97 ± 0.24 80.95 ± 0.25 80.14 ± 0.12 72.08 ± 0.17
E1 Base FT Fine-tune all layers 79.14 ± 0.22 82.12 ± 0.02 70.67 ± 1.64 69.04 ± 1.29 9.73 0.12
E2 Large FT Fine-tune all layers 82.76 ± 0.57 86.22 ± 0.44 77.33 ± 1.46 71.77 ± 1.22 53.26 0.65
E3 4o-mini Base FT Prompt, Label 82.50 ± 0.34 86.40 ± 0.15 75.33 ± 1.22 70.88 ± 1.20 9.73 0.12
E4 4o-mini Large FT Prompt, Label 83.80 ± 0.43 87.71 ± 0.27 78.73 ± 1.12 71.77 ± 1.10 53.26 0.64
E5 4o-mini Large FT Prompt, Label, Examples (Few shot) 83.42 ± 0.30 86.94 ± 0.28 79.50 ± 1.12 72.33 ± 0.49 53.26 0.64
E6 4o-mini Large FT Prompt, Label, Balanced Ex. (Few shot) 82.98 ± 0.42 86.28 ± 0.62 79.87 ± 0.36 71.98 ± 0.75 53.26 0.64
E7 4o-mini Large FT Prompt, Probs 83.27 ± 0.37 86.60 ± 0.23 79.41 ± 0.69 72.26 ± 1.03 53.26 0.64
E8 4o-mini Large FT Prompt, Label, Probs 83.66 ± 0.32 87.22 ± 0.28 79.98 ± 0.36 71.78 ± 1.06 53.26 0.64
E9 4o-mini Large FT Prompt, Label, Probs, Examples 83.19 ± 0.39 86.58 ± 0.60 78.99 ± 0.42 71.94 ± 0.64 53.26 0.64
E10 4o-mini FT Fine-tune w/prompt 86.70 ± 0.11 89.65 ± 0.30 87.00 ± 0.13 75.83 ± 0.21 33.15 0.38
E11 4o-mini FT 5 Fine-tune w/prompt (5 epochs) 84.86 ± 0.13 87.74 ± 0.13 86.22 ± 0.40 75.38 ± 0.32 165.75 1.95
E12 4o-mini FT-M Minimal fine-tune 86.51 ± 0.06 89.57 ± 0.18 87.13 ± 0.22 75.74 ± 0.17 16.60 0.19
E13 4o-mini FT Base FT Prompt, Label, FT w/prompt 81.06 ± 0.52 84.67 ± 0.14 73.06 ± 2.03 69.70 ± 1.77 42.88 0.53
E14 4o-mini FT-L Base FT Prompt, Label, FT w/prompt, label 81.84 ± 0.26 85.20 ± 0.06 77.29 ± 1.22 70.70 ± 1.39 49.31 0.60
E15 4o-mini FT-L 5 Base FT Prompt, Label, FT w/prompt, label (5 epochs) 83.67 ± 0.30 86.38 ± 0.51 81.19 ± 0.62 75.02 ± 0.03 207.64 2.48
E16 4o-mini FT Large FT Fine-tune w/prompt 83.94 ± 0.09 87.57 ± 0.11 80.17 ± 0.28 72.46 ± 0.00 86.41 1.03
E17 4o-mini FT-L Large FT Fine-tune w/prompt, label 84.12 ± 0.06 87.58 ± 0.10 80.75 ± 0.20 73.34 ± 0.06 92.84 1.10
E18 4o-mini FT-L 5 Large FT Fine-tune w/prompt, label (5 epochs) 84.83 ± 0.06 87.75 ± 0.17 84.37 ± 0.81 75.59 ± 0.01 251.17 2.96
E19 4o Large FT Prompt, Label 83.19 ± 0.01 85.71 ± 0.00 82.06 ± 0.11 73.48 ± 0.06 53.26 0.64
E20 4o Large FT Prompt, Label, Examples (Few shot) 83.29 ± 0.28 86.11 ± 0.14 81.48 ± 0.07 72.96 ± 1.27 53.26 0.64
E21 4o Large FT Prompt, Label, Balanced Ex. (Few shot) 83.19 ± 0.28 86.01 ± 0.19 81.04 ± 0.21 72.88 ± 1.03 53.26 0.64
E22 4o Large FT Prompt, Probs 82.99 ± 0.47 86.37 ± 0.45 78.42 ± 1.05 71.90 ± 1.04 53.26 0.64
E23 4o Large FT Prompt, Label, Probs 83.31 ± 0.33 86.69 ± 0.35 79.46 ± 0.33 72.17 ± 0.97 53.26 0.64
E24 4o Large FT Prompt, Label, Probs, Examples 83.04 ± 0.42 86.53 ± 0.29 78.47 ± 1.00 71.83 ± 1.20 53.26 0.64
E25 4o FT Fine-tune w/prompt 86.79 ± 0.06 90.46 ± 0.03 88.14 ± 0.28 73.09 ± 0.01 276.24 3.18
E26 4o FT-M Minimal fine-tune 86.99 ± 0.00 90.57 ± 0.00 89.00 ± 0.00 73.99 ± 0.00 138.37 1.59
E27 4o FT Large FT Fine-tune w/prompt 84.03 ± 0.30 87.90 ± 0.13 80.01 ± 0.73 72.00 ± 1.15 329.50 3.93
E28 4o FT-L Large FT Fine-tune w/prompt, label 84.37 ± 0.19 87.81 ± 0.09 81.28 ± 1.03 73.10 ± 0.66 383.10 4.55
  • Bold = best overall, highlighted = best in section
  • 1 Some ID numbers changed from their original ID in the research repo.
  • 2 GPT fine-tuning types: FT = fine-tune with prompt, FT-M = minimal without prompt, FT-L = with prompt including ELECTRA label, FT 5 = 5 epochs vs. 1
  • 3 Merged dataset: Combination of test splits from DynaSent R1/R2 and SST-3
  • 4 Each experiment was run twice with different random seeds (42, 123) and temperature (0.0, 0.1); values reported are means ± standard deviations. Standard deviations are based on two runs (n=2) and should be interpreted with caution.
  • 5 Cost: FT = Fine-tuning cost, no inference-time API charges. Ratio is FT cost divided by F1 score.
Our experiments revealed significant differences in performance across baseline, fine-tuning, and collaborative scenarios. See Table 6 for the mean macro average F1 between the two runs. Appendix A has the raw data of each run.
Baselines. Regarding baselines, both GPT models outperformed the ELECTRA classifiers, with GPT-4o achieving a 79.97 mean macro F1 and GPT-4o-mini scoring 79.41, compared to ELECTRA Base (69.51) and Large (67.94). This demonstrates the strong zero-shot capabilities of GPT models.
Fine-tuning. Fine-tuning improved performance across all models. ELECTRA Base’s mean macro F1 increased from 69.51 to 79.14, while ELECTRA Large showed greater gains, improving from 67.94 to 82.76. This improvement is the result of fine-tuning all layers—the baselines had the same classifier head. Fine-tuned GPT models had the highest scores (see Figure 2), with GPT-4o-mini FT rising from 79.41 to 86.70, and GPT-4o FT-M achieving 86.99 with the minimal template.
Figure 1: Change in Mean F1 from Adding Predictions
Sharing Predictions. The effect of adding ELECTRA predictions to GPT prompts depended on if the GPT model was fine-tuned (see Figure 1 for the differences in mean F1). Sharing ELECTRA Base predictions with GPT-4o-mini (not fine-tuned) significantly improved the macro F1 in round one from 79.52 to 82.74 (p < 0.0001, McNemar’s test and bootstrap analysis), a +3.22 gain. There was an even greater gain of +3.97 points when ELECTRA Large predictions were shared (from 79.52 to 83.49, p < 0.0001). Similarly, including ELECTRA Large predictions with GPT-4o improved the macro F1 from 80.14 to 83.18 (p < 0.001) in round one, a +3.04 gain.
Figure 2: Mean Macro F1 Scores on Merged Dataset by Experiment
Figure 3: Impact of Context on Mean F1 Score
However, sharing ELECTRA predictions with fine-tuned GPT models actually decreased performance. GPT-4o-mini FT’s mean macro F1 dropped from 86.70 to 81.06 when including ELECTRA Base predictions, and to 81.84 when fine-tuned with the predictions included in the prompt. Similarly, GPT-4o FT’s mean F1 fell from 86.79 to 84.03 when including ELECTRA Large predictions, and to 84.37 when fine-tuned with them.
Few-shot Examples. Some contexts performed better than others for specific model combinations (see Figure 3). Providing few-shot examples in addition to the predicted label was mostly the same or worse than using the label alone. However, when looking at the more challenging DynaSent Round 2 dataset, GPT-4o-mini saw some benefit. Including just the ELECTRA Large predicted label produced a mean macro F1 of 78.73. Adding examples increased the mean macro F1 to 79.50 (+0.77), and balanced examples increased it to 79.87 (+1.14).
Sharing Probabilities. Using probabilities instead of (or in addition to) the predicted label was mostly the same or worse than using the label alone. However, similar to using examples, the more challenging datasets saw some benefit. For DynaSent Round 2, GPT-4o-mini had a mean macro F1 of 78.73 with just the ELECTRA Large predicted label. Using probabilities instead changed it to 79.41, and using the label with probabilities increased it to 79.98. A similar minor improvement was seen with SST on this dataset.
Datasets. Performance also varied across datasets. GPT-4o FT-M achieved the top scores on DynaSent R1 (90.57 mean macro F1) and DynaSent R2 (89.00). Surprisingly, GPT-4o-mini FT—the smaller model—performed best on SST-3 with a 75.83 mean macro F1, exceeding even GPT-4o FT’s performance of 73.99.
Figure 4: Incorrect Predictions by Label (Round 1)
7 Analysis
However, following ELECTRA’s predictions had mixed results. When GPT-4o-mini changed its decision and followed ELECTRA Base FT in round one, it was correct 548 times and wrong 412 times (+136 net improvement, 57.08% success rate). When GPT-4o changed its decision and followed ELECTRA Large FT, it was correct 521 times and wrong 481 times (+40 net improvement, 52% success rate).
H2. Improvement would be less for fine-tuned GPTs. H2 was supported more strongly than anticipated. For a fine-tuned GPT model, including the ELECTRA prediction actually decreased performance (see Figure 1). Initially, we thought this was because the fine-tuning context did not include the ELECTRA prediction in the prompt. But we still saw a decrease in performance (although less) when it was included.
GPT was trained to be a helpful chat assistant that thinks through problems, evaluates information critically, and synthesizes knowledge. When presented with an ELECTRA prediction, it can decide when to follow the prediction and when to rely on its own judgment. In contrast, when fine-tuned specifically for sentiment classification, the model is optimized to map directly from input text to sentiment labels, potentially bypassing the critical thinking that made the base model effective at filtering ELECTRA’s suggestions.
It may also be that it takes more time to learn the nuances of when ELECTRA is right vs. wrong. When we extended fine-tuning from 1 to 5 epochs, GPT was better able to discriminate ELECTRA’s predictions. The follow rate decreased from 96.34% to 89.66%, and the discrimination gap—the difference between following correct and incorrect predictions—went from 16.89% to 35.95%. Although additional epochs improved performance, it comes with a significant increase to the cost: \$207.64 to \$251.17 vs. only \$33.15 to just fine-tune GPT-4o-mini—with better performance.
H3. Format of prediction would impact performance. Although using the predicted label alone was best, GPT-4o-mini saw more variability in performance between the different contexts, as shown in Figure 3. However, for GPT-4o, there was not much of a difference when looking at the mean change in F1 score.
H4. Including examples would improve performance. Contrary to this hypothesis, including examples did not improve performance for either GPT-4o-mini or GPT-4o when looking at the mean change in F1 score. It either had no effect (GPT-4o) or actually decreased performance (GPT-4o-mini).
8 Conclusion
This research investigated collaborative approaches to sentiment classification between bidirectional transformers and LLMs. Our results show that augmenting prompts with predictions from a fine-tuned ELECTRA can significantly improve performance when the GPT model is not fine-tuned—up to +4.39 points of gain in the mean macro F1 score. Including probabilities or similar examples improved performance slightly on the more challenging datasets. However, this collaborative benefit disappeared when the GPT models were fine-tuned.
It is possible that the fine-tuning process hinders the critical thinking that’s present in the default GPT models. Fine-tuning over more epochs improves the situation—GPT gets smarter about when to follow ELECTRA vs. when to ignore it. However, this comes at a significant cost increase—six to seven times the cost of fine-tuning GPT-4o-mini alone, and the performance is still not comparable.
Our findings offer several cost-effective paths for sentiment analysis projects. For organizations that can fine-tune via API, GPT-4o-mini FT offers nearly equivalent performance to GPT-4o FT-M (86.70 vs 86.99 mean macro F1) at 76% lower cost (\$0.38 vs \$1.59/F1 point). For those with data privacy concerns or resource constraints, GPT-4o-mini with ELECTRA Base FT had the best cost/performance ratio (\$0.12/F1 point). Projects that need to stay completely local can fine-tune ELECTRA Large, which outperformed both base GPT models.
9 Limitations
The cost/performance evaluation only considered the fine-tuning costs to achieve the reported macro F1 on the test set. In practice, there may be ongoing costs for inference time API calls for hosted GPT models. The time and cost required to fine-tune different datasets will vary, and this will affect the relative cost/performance differences.
Our research was limited to GPT-4o and GPT-4o-mini from OpenAI, but there are many GPT-style models available. Some of these are open source and can be fine-tuned locally or on hosted compute resources, such as Llama 2 or 3 from Meta (Touvron et al., 2023; Dubey et al., 2024). The cost of fine-tuning an open source GPT model without API fees was not evaluated.
Many experimental runs involved prompting the GPT models, and these prompts could be further optimized with frameworks like DSPy. Some of the prompts are verbose, and these additional tokens could dilute the signal of the ELECTRA context.
Although we quantified changes in GPT classification decisions before and after ELECTRA predictions were included, we did not have time to thoroughly explore why the ELECTRA prediction was ignored in some cases, but followed in others. Future research could prompt the model to explain their decision, and the responses could be analyzed. This additional instruction would likely affect the classification decision—similar to Chain of Thought (Wei et al., 2022)—and would be a new experimental scenario.
We tried to give the GPT models a chance to learn when ELECTRA was right or wrong during fine-tuning (GPT-4o FT-L, GPT-4o-mini FT-L). There may be a more direct way of fine-tuning GPT so that it learns when to trust ELECTRA’s prediction, and when to decide for itself.
The data used for this research was a merge of movie and business reviews from SST-3 and DynaSent R1/R2. The majority of the reviews were written in English and relatively short in length. Assessing generalization to other domains, languages and longer lengths would require future research.
10 Acknowledgments
Thank you to the Stanford Engineering Center for Global and Online Education (CGOE) for the opportunity to conduct this research as part of the Natural Language Understanding class (XCS224U), and to Professor Christopher Potts, Insop Song, and Petra Parikova for your support and guidance.
Appendix A Results of Individual Runs
Table 7: Round 1 Results
Merged DynaSent R1 DynaSent R2 SST-3 Cost (\$)